### v1

# Gold after Randomized Sand: Model-X Split Knockoffs for Controlled Transformation Selection

Abstract
Controlling the False Discovery Rate (FDR) in variable selection is crucial for reproducibility and preventing over-selection, particularly with the increasing prevalence of predictive modeling. The Split Knockoff method, a recent extension of the canonical Knockoffs framework, offers finite-sample FDR control for selecting sparse transformations, finding applications across signal processing, economics, information technology, and the life sciences. However, its current formulation is limited to fixed design settings, restricting its use to linear models. The question of whether it can be generalized to random designs, thereby accommodating a broader range of models beyond the linear case – similar to the Model-X Knockoff framework – remains unanswered. A major challenge in addressing transformational sparsity within random design settings lies in reconciling the combination of a random design with a deterministic transformation. To overcome this limitation, we propose the Model-X Split Knockoff method. Our method achieves FDR control for transformation selection in random designs, bridging the gap between existing approaches. This is accomplished by introducing an auxiliary randomized design that interacts with both the existing random design and the deterministic transformation, enabling the construction of Model-X Split Knockoffs. Like the classical Model-X framework, our method provides provable finite-sample FDR control under known or accurately estimated covariate distributions, regardless of the conditional distribution of the response. Importantly, it guarantees at least the same selection power as Model-X Knockoffs when both are applicable. Empirical studies, including simulations and real-world applications to Alzheimer’s disease imaging and university ranking analysis, demonstrate robust FDR control and suggest improved selection power over the original Model-X approach.
1 Introduction
Variable selection is a fundamental problem in statistical research, focusing on identifying which variables from a large candidate pool significantly influence an outcome. A key challenge in this area is ensuring reproducibility and replicability, as the inclusion of more predictors often increases predictive power but risks over-selection in finite samples. To address this, controlling the False Discovery Rate (FDR) has become a central concern. The seminal work of Benjamini and Hochberg (1995) introduced the concept of FDR and proposed a method to control it in settings with independent hypothesis tests or orthogonal designs. This groundbreaking contribution laid the foundation for extensive theoretical and practical advancements in FDR control (Benjamini and Yekutieli, 2001, Storey, 2002, Reiner et al., 2003, Benjamini, 2010).
Building on this, Barber and Candès (2015) introduced the Knockoff method, which extended FDR control with finite-sample guarantees to general fixed design settings, including sparse linear models with arbitrary design matrices. This represented a significant advance over earlier methods, which were often limited to specific design classes. The Knockoff framework has since inspired considerable research (Dai and Barber, 2016, Xu et al., 2016, Barber and Candès, 2019).
Further generalization was recently achieved by Cao et al. (2024a, b), who introduced the concept of transformational sparsity. While traditional Knockoff methods focus on sparsity in the original measured covariates, transformational sparsity considers sparsity in predetermined transformations (often linear) of the parameters. This paradigm is particularly relevant in applications where sparsity manifests in derived features rather than raw variables. For example, in genomic research, disease mechanisms often involve biological pathways (linear combinations of gene effects) rather than isolated genes (Ramanan et al., 2012); in neuroimaging, wavelet transforms are applied to MRI data to extract meaningful patterns (Donoho and Johnstone, 1995, Bullmore et al., 2004); and in economics, temporal contrasts like trend filtering are often analyzed instead of raw features (Kim et al., 2009). To address this, Cao et al. (2024a) introduced the Split Knockoff framework, which extends the original Knockoff method to control the false discovery rate (FDR) with finite-sample guarantees under transformational sparsity in fixed design settings. This framework was further generalized to control directional FDR in Cao et al. (2024b), which finds successful applications in Alzheimer’s disease image analysis and statistical ranking with multiple comparisons.
In another direction, Candes et al. (2018) introduced the Model-X Knockoff framework, which enables FDR control in random design settings. This method leverages knowledge of the marginal distribution of the design matrix and identifies the Markov blanket – the smallest subset of variables that renders the outcome independent of all others – allowing variable selection in a wide range of statistical models. The Model-X framework represents a significant advance, accommodating diverse modeling scenarios and inspiring substantial research (Barber et al., 2020, Romano et al., 2020, Ren and Candès, 2023, Ren and Barber, 2024).
However, the Model-X framework is not directly compatible with transformational sparsity, where sparsity is defined with respect to deterministic transformations of the parameters rather than the associated variables themselves. For instance, in pairwise comparison problems, identifying that an outcome depends on two variables (as part of the Markov blanket) does not indicate whether these variables exert the same or different effects on the outcome. While Model-X Knockoffs identify influential variables, transformational sparsity concerns specific relationships or null hypotheses about transformed parameters, such as testing the equality of effects.
To address this gap, this paper introduces the Model-X Split Knockoff method, which extends the Model-X Knockoff framework to handle transformational sparsity in random design settings. This novel method provides FDR control with finite-sample guarantees under transformational sparsity across a broad range of statistical models, beyond the linear settings considered in earlier works. Moreover, as will be demonstrated, the Model-X Split Knockoff method achieves at least the same selection power as the original Model-X Knockoff method and often outperforms it in empirical applications when both methods are applicable.
Specifically, this paper considers models where the conditional expectation of the response $Y\in\mathbb{R}$ given the random designs $X:=(X_{1},\dots,X_{p})\in\mathbb{R}^{p}$ subject to the marginal distribution $P^{*}_{X}$, as well as the deterministic transformation of parameters, is defined as follows:
$\mathbb{E}(Y|X)=f(X\beta^{*}),\quad\gamma^{*}=D\beta^{*},$ (1)
where $f:\mathbb{R}\to\mathbb{R}$ is the link function, and $D\in\mathbb{R}^{m\times p}$ is a predetermined linear transformation matrix that induces sparsity in the transformed coefficient vector $\gamma^{*}\in\mathbb{R}^{m}$. Transformation selection focuses on identifying the support set of $\gamma^{*}$, i.e. $\mathcal{H}_{1}^{\gamma}=\{i:\gamma^{*}_{i}\neq 0\}$, rather than directly identifying the support of $\beta^{*}\in\mathbb{R}^{p}$. For example, in multiple comparison problems, $D$ is the graph difference matrix such that $(D\beta^{*})(i,j):=\beta^{*}_{i}-\beta^{*}_{j}$ for each pair $(i,j)$ in an edge set. In this case, $\gamma^{*}_{i}=0$ corresponds to the null hypothesis $\mathcal{H}^{i,j}:\beta_{i}^{*}=\beta_{j}^{*}$, meaning that covariate $X_{i}$ has the same conditional effect as $X_{j}$.
A significant challenge in working with model (1) arises from its inherent structure, which combines the random design $X$ with a deterministic transformation $D$, ultimately defining $\gamma^{*}$, whose sparsity is the primary focus. In this transformational sparsity scenario, there is no straightforward design matrix for $\gamma^{*}$ to directly work with, as neither $X$ alone nor $D$ directly maps $\gamma^{*}$ to the observed outcome $Y$. Instead, the randomness of $X$ becomes entangled with the deterministic transformation $D$, resulting in an implicit subspace constraint. This unique scenario poses challenges that lie outside the scope of the traditional Model-X framework, introducing substantial complexities for both methodological development and theoretical analysis.
To address these challenges, we adopt the Split Knockoffs approach introduced by Cao et al. (2024a), which considers the lifted parameter space $(\beta,\gamma)$ by relaxing the subspace constraint in (1) to its Euclidean neighborhood. To reconcile the random design $X$ with the deterministic design implicitly imposed by the subspace constraints, we introduce an auxiliary random vector $A\in\mathbb{R}^{m}$. This random vector is specifically constructed to map the constraint on $\gamma^{*}$ and its neighborhood to the measurement associated with the response, effectively introducing a randomized design associated with $\gamma^{*}$ into the model. This facilitates the construction of Model-X Split Knockoffs for inference on the transformed parameters $\gamma^{*}$. Specifically, the auxiliary randomized design $A$ transforms $X\beta^{*}$ into the following random design on the lifted parameter $(\beta^{*},\gamma^{*})$:
$X\beta^{*}=\widetilde{X}^{A}\beta^{*}+A\gamma^{*},\quad\widetilde{X}^{A}:=X-AD.$ (2)
This decomposition holds because $A\gamma^{*}=AD\beta^{*}$ for all $A$. In this formulation, $\beta^{*}$ associated with the transformed random design $\widetilde{X}^{A}=X-AD$, while $\gamma^{*}$ — which captures the transformational sparsity — is linked to the auxiliary randomized design $A$. This decomposition enables us to design Model-X split knockoffs for $A$ conditional on $\widetilde{X}^{A}$, provided some exchangeability and independence conditions are satisfied (see Section 2.2 for details). Provable false discovery rate (FDR) control is guaranteed as long as $P^{*}_{X}$ is known, bootstrapped, or can be accurately estimated (see Section 3). Furthermore, the randomized design $A$ and the relaxation of the subspace constraint, from $\gamma^{*}=D\beta^{*}$ to its Euclidean neighborhood, can be leveraged to enhance selection power (see Section 2.3 and discussions therein).
Since the proposed method is grounded in the Model-X framework and does not require specifying the conditional distribution $P_{Y|X}$, it accommodates more complex models beyond simple linear settings. Additionally, this separation provides greater flexibility to optimize selection power compared to the original Model-X Knockoff method when $D=I$, as it imposes no assumptions on the distribution of $A$. Benefiting from this flexibility, the Model-X Split Knockoff method is guaranteed to achieve at least the same power as the Model-X Knockoff method, and empirical results demonstrate that it often achieves superior selection power in practice.
While the choice of the auxiliary random vector $A$ could, in principle, be arbitrary, subsequent steps in the Model-X Split Knockoff procedure are substantially dependent on the joint distribution of $(X,A)$. Although a general algorithm that can operate with an arbitrary joint distribution of $(X,A)$ exists, this paper will introduce specific constructions of $A$ tailored to canonical examples, such as pairwise comparisons. The motivation for developing these tailored constructions is to facilitate more efficient computation and to improve statistical power, particularly in scenarios where the distribution of $X$ is unknown and must be estimated from data.
To evaluate the performance of our proposed method, we conduct extensive experiments on both simulated and real-world datasets, including data for Alzheimer’s Disease and world college rankings. Simulation experiments using both regression and pairwise comparison data from random graphs validate the method’s effectiveness. For the Alzheimer’s Disease dataset, the method successfully identifies abnormal regions and connections associated with the disease. Furthermore, it identifies important and significant gaps between college pairs in the world college ranking dataset.
The paper is organized as follows: Section 2 introduces the methodology of Model-X Split Knockoff, including specific construction methods for canonical examples. The theoretical results for FDR control of the proposed method are presented in Section 3. Simulation experiments are detailed in Section 4, while experiments on real-world datasets are discussed in Section 5 and Section 6.
2 Methodology
In this paper, we address the transformational sparsity problem, modeled in Equation (1) and quoted here for convenience:
$\mathbb{E}(Y|X)=f(X\beta^{*}),\quad\gamma^{*}=D\beta^{*}.$
The primary goal of this paper is to estimate the support set of $\gamma^{*}$, defined as $\mathcal{H}^{\gamma}_{1}=\{i:\gamma^{*}_{i}\neq 0\}$, while controlling the false discovery rate. A key challenge arises from the interplay of a random design $X$ and a deterministic transformation $D$, where neither directly maps $\gamma^{*}$ to the outcome $Y$.
To overcome this, we draw inspiration from the variable splitting technique recently proposed by Cao et al. (2024a). We introduce a new auxiliary design for $\gamma^{*}$and integrate it into the model to facilitate inference on $\gamma^{*}$. The specifics of constructing this valid design are detailed below.
2.1 Variable Splitting and Auxiliary Randomized Design
An auxiliary random vector $A=(A_{1},\dots,A_{m})$ is generated as in (2) to manually construct the design for $\gamma^{*}$, enabling the reformulation of model (1) as:
$\mathbb{E}(Y|\widetilde{X}^{A},A)=f(\widetilde{X}^{A}\beta^{*}+A\gamma^{*}).$ (3)
Recall that the equation holds because $\widetilde{X}^{A}:=X-AD$ and $A\gamma^{*}=D\beta^{*}$.
In this way, we successfully inserted an auxiliary $A$ for the parameter of interest $\gamma^{*}$, and lifted the model (1) into a higher dimensional space including both $\beta$ and $\gamma$. With Equation (3), a Split Knockoff copy $\widetilde{A}$ following certain necessary constraints could be constructed to enable variable selection with the false discovery rate control on $\gamma^{*}$. After such construction, standard procedures of Split Knockoffs as described in Cao et al. (2024a) could be applied to finalize the procedure. In the following section, we will proceed to the step of constructing the Split Knockoff copy.
2.2 Construction of Model-X Split Knockoffs
Similarly to (Candes et al., 2018, Cao et al., 2024a), the Model-X Split Knockoff requires the construction of so-called Split Knockoff copies to achieve the false discovery rate control which will be shown in Section 3. The random vector $\widetilde{A}=(\widetilde{A}_{1},\dots,\widetilde{A}_{m})$ is called a Split Knockoff copy with respect to $(\widetilde{X}^{A},A)$ if the following conditions are satisfied:
  • Exchangeability: for any subset $S\subseteq\{1,\dots,p\}$,
    $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}\overset{\mathcal{D}}{=}(\widetilde{X}^{A},A,\widetilde{A}).$ (4)
  • Conditional independence: $\widetilde{A}\perp\!\!\!\perp Y|(\widetilde{X}^{A},A)$.
Here, $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}$ means swap features in $S$ between $A$ and $\widetilde{A}$. For example, if $A\in R^{2},S=\{1\}$, then
$(\widetilde{X}^{A},A_{1},A_{2},\widetilde{A}_{1},\widetilde{A}_{2})_{\text{swap}(S)}=(\widetilde{X}^{A},\widetilde{A}_{1},A_{2},A_{1},\widetilde{A}_{2}).$
In the cases where the distribution of $X$ is well known, for any choice of $A$, the $\widetilde{A}$ satisfying the exchangeability condition (4) can be constructed in a sequential way as shown in Algorithm 1, similar to Candes et al. (2018).
Algorithm 1 Sequential Generation of Split Knockoff Copies
the distribution of $X$
Generate random variable $A$ and the respective random variable $\widetilde{X}^{A}$
for $j=1:m$ do
     sample $\widetilde{A}_{j}$ from the conditional distribution: $\mathcal{L}(A_{j}|A_{-j},\widetilde{A}_{1:j-1},\widetilde{X}^{A})$
end for
The proof of Algorithm 1 satisfies the exchangeability is given in Supplementary Material Section A.
Although Algorithm 1 gives the general construction of the Split Knockoff copy $\widetilde{A}$, it might be facing computational challenges in practice as the computational cost enlarges quickly as $m$ enlarges. Moreover, the generated Split Knockoff copies can be highly correlated which harms the selection power. In the following, we deal with two canonical examples to present refined constructions of Split Knockoff copies, which leads to lower computational costs and better selection power.
2.2.1 Construction for Normal Distributions
In the cases where the distribution of $X$ follows or can be well-approximated by the normal distribution $\mathcal{N}(0,\Sigma_{X})$, the nature choice of the pair $(A,\widetilde{A})$ is to generate it from a joint normal distribution of $(X,A,\widetilde{A})$, as suggested by Candes et al. (2018). Specially, as we shall shown later, among all choices of joint normal distributions, the choices which makes the “intercept” $\widetilde{X}^{A}$ independent from the auxiliary design $A$ are expected to achieve higher selection power.
The generation process begins with $A$ being drawn from
$(X,A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{X}&\Sigma_{AX}^{T}\\ \Sigma_{AX}&\Sigma_{A}\end{array}\right)\right\},$ (9)
where $\Sigma_{A}$ and $\Sigma_{AX}$ are two suitable matrices which could be used to optimize the selection power. Then by definition of $\widetilde{X}^{A}=X-AD$, there holds
$(\widetilde{X}^{A},A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{\widetilde{X}^{A}}&\Sigma_{AX}^{T}-D^{T}\Sigma_{A}\\ \Sigma_{AX}-\Sigma_{A}D&\Sigma_{A}\end{array}\right)\right\}.$ (14)
where $\Sigma_{\widetilde{X}^{A}}=\Sigma_{X}-\Sigma_{XA}^{T}D-\Sigma_{XA}D^{T}+D^{T}\Sigma_{A}D$. Then the conditional distribution of $A|\widetilde{X}^{A}$ evaluated at ${\widetilde{X}^{A}}=\widetilde{x}^{A}$ in Algorithm 1 is
$A|(\widetilde{X}^{A}=\widetilde{x}^{A})\sim\mathcal{N}\left\{(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T},\Sigma_{A}-(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\Sigma_{AX}^{T}-D^{T}\Sigma_{A})\right\}$ (15)
Then conditional on the observation $\widetilde{X}^{A}=\widetilde{x}^{A}$, the Split Knockoff copy $\widetilde{A}$ could be generated from the joint distribution
$(A,\widetilde{A})|(\widetilde{X}^{A}=\widetilde{x}^{A})\sim\mathcal{N}\left\{\left(\begin{array}[]{c}(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T}\\ (\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T}\end{array}\right),\left(\begin{array}[]{cc}\widetilde{\Sigma}_{A}&\widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})\\ \widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})&\widetilde{\Sigma}_{A}\end{array}\right)\right\},$ (20)
where $\widetilde{\Sigma}_{A}=\Sigma_{A}-(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\Sigma_{AX}^{T}-D^{T}\Sigma_{A})$, $s\in\mathbb{R}^{m}$ is an non-negative vector which makes the above covariance matrix semi-positive definite. The proof of the argument that this construction satisfies the exchangeability can be directly obtained from the normal distributions and will be omitted.
Now we are going to specify the choice of $\Sigma_{A}$ and $\Sigma_{AX}$ in order to optimize the selection power.
  1. 1
    Choice of $\Sigma_{AX}$. As discussed in Barber and Candès (2015), Candes et al. (2018), the non-negative vector $s$ in Equation (20) should be as large as possible to maximize the selection power. It could be verified that $s$ needs to satisfy
    $s\preceq 2\widetilde{\Sigma}_{A}\preceq 2\Sigma_{A},$
    where for the second inequality, the equation holds if and only if $\Sigma_{AX}=\Sigma_{A}D$. This gives us evidence that taking $\Sigma_{AX}=\Sigma_{A}D$, i.e. $\widetilde{X}^{A}$ independent from $A$ in (14), helps optimize the selection power.
  2. 2
    Choice of $\Sigma_{A}$. Another intuition on improving the selection power is to improve the incoherence conditions as mentioned in Cao et al. (2024a). In this spirit, one shall take $\Sigma_{A}=\alpha I_{m}$ for some $\alpha>0$ to improve the selection power.
In summary, the choice of $\Sigma_{A}$ and $\Sigma_{AX}$ to optimize the selection power is $\Sigma_{A}=\alpha I_{m}$ and $\Sigma_{AX}=\alpha D$. In this case, the joint distribution over $(X,A)$ is
$(X,A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{X}&\alpha D^{\top}\\ \alpha D&\alpha I_{m}\end{array}\right)\right\}.$ (25)
where $\alpha$ satisfies $\Sigma_{X}-\alpha D^{T}D\succeq 0$. In this setting, $\widetilde{A}$ is generated from
$(A,\widetilde{A})\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{m}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\alpha I_{m}&\alpha I_{m}-\mathrm{diag}(\mathbf{s})\\ \alpha I_{m}-\mathrm{diag}(\mathbf{s})&\alpha I_{m}\end{array}\right)\right\}.$ (30)
This specific choice of $A$ and $\widetilde{A}$ generates independent designs for features of interests, which enhances the selection power as validated by experiments in Section 4.1. Moreover, when the covariance matrix $\Sigma_{X}$ is unknown, one can see a good estimation of such covariance matrices leads to a robust FDR control in Section 3.2.
2.2.2 Construction for Pairwise Comparisons
In this section, we discuss the case of pairwise comparisons where the distribution for the design matrix $X$ can not be approximated by Gaussian distributions.
For example, let a collection of pairwise comparison be represented by $\mathbf{X}\in\mathbb{R}^{n\times p},\mathbf{Y}\in\mathbb{R}^{n}$ where each row $X\in\mathbb{R}^{p}$ contains a random paired comparison such that $1$ denotes the first term in the random pair and $-1$ denotes the second term in the random pair and other values are 0. For example, $X=(1,-1,0,\dots,0)$ means that the comparison is done between the first and second objects. Each element $Y\in\mathbb{R}$ of $\mathbf{Y}$ is modeled from
$\mathbb{P}(Y=1|X)=f(X\beta^{*}),\ \mathbb{P}(Y=-1|X)=f(-X\beta^{*}),$ (31)
where $\beta^{*}\in\mathbb{R}^{p}$ represents the unknown true scores for $p$ objects, and $f(x)$ is a cumulative distribution function which is symmetrical at 0, i.e. it satisfies $f(x)=1-f(-x)$. In order to conduct pairwise comparisons, let $D$ be the graph difference operator of a fully connected graph. [^1] The construction detailed in this section is applicable when $D$ contains all the edges present in $\mathbf{X}$. Should this condition not be met, one can simply augment $D$ to include all required edges. Then the pairs with large differences will be selected in the model (1).
Sequential Constructions
One direct way to construct the pair $(A,\widetilde{A})$ is the sequential construction following Algorithm 1 with certain choice of $A$. The detailed construction for the above pairwise comparisons setting is shown in Supplementary Material Section F. However, as will be demonstrated later in Section 4.2, these constructions not only incur high computational costs but also exhibit reduced selection power due to the inherent dependency among $(\widetilde{X}^{A},A,\widetilde{A})$. This motivates our search for a simultaneous approach to construct Split Knockoff copies. Such an approach would not only reduce computational expense but also enhance selection power by ensuring $\widetilde{X}^{A}$ is independent of $A$, a characteristic previously demonstrated to be advantageous for selection power in the context of normal distributions.
Bootstrap+ Constructions
In this section, we present a novel method for generating Split Knockoff copies, which we term the bootstrap+ construction. This approach simultaneously constructs the pair $(A,\widetilde{A})$ and incorporates bootstrap sampling of the design matrix $\mathbf{X}^{r}$ after augmenting it with rows of $0_{p}$.
This construction initiates with an augmented bootstrap resampling procedure. Each row $(X^{r},Y^{r})$ of the resulting augmented and resampled design matrix $\mathbf{X}^{r}$ and response vector $\mathbf{Y}^{r}$ is generated from the distribution of $X$ by:
  1. 1.
    $\mathbb{P}(X^{r}=0_{p})=\mathbb{P}(X^{r}\neq 0_{p})=\frac{1}{2}$.  [^2] One can not use a general probability $\mathbb{P}(X^{r}=0_{p})=p_{0}\neq 1/2$ here since $p_{0}\neq 1/2$ will destroy the exchangeability of $\tilde{A}$ and $A$.
  2. 2.
    When $X^{r}=0_{p}$, $Y^{r}$ is generated through (31), i.e. $\mathbb{P}(Y^{r}=1)=\mathbb{P}(Y^{r}=-1)=\frac{1}{2}$.
  3. 3.
    When $X^{r}\neq 0_{p}$, $X^{r}$ takes a random row from $\mathbf{X}$, and $Y^{r}$ takes the corresponding row in $\mathbf{Y}$.
This resampling enables the following construction of the Split Knockoff copies based on $(X^{r},Y^{r})$ at a much lower computational cost.
  1. 1.
    The random vector $A\in\mathbb{R}^{m}$ is generated to satisfy $\widetilde{X}^{A}_{r}:=X^{r}-AD=0_{p}$, i.e. if $X^{r}=0$ then $A=0$; if $X^{r}$ takes a random row of $\mathbf{X}$ then $A$ selects the same row in $D$. In this way, $\widetilde{X}^{A}_{r}$ is clearly independent from $A$, which should help improve the selection power as shown in the previous Gaussian case of $X\sim\mathcal{N}(0,\Sigma_{X})$.
  2. 2.
    The random vector $\widetilde{A}$ is generated based on $A$ by
    $\mathbb{P}(\widetilde{A}=0_{m}|A\neq 0_{m})=1,$
    $\mathbb{P}(\widetilde{A}=e_{i}|A=0_{m})=\mathbb{P}(A=e_{i}|A\neq 0_{m}),$
    for any canonical basis vector $e_{i}$.
With the above construction, one can easily verify that $\widetilde{X}^{A}_{r}\equiv 0_{p}$, while the vector $[A,\widetilde{A}]\in\mathbb{R}^{2m}$ is the one-hot vector with one element being 1 and other elements being zeros. Moreover, $\mathbb{P}[A_{j}=1]=\mathbb{P}[\widetilde{A}_{j}=1]$ for all $j$, which directly leads to the exchangeability.
This procedure clearly lets $\widetilde{X}^{A}_{r}$ independent from $A$, which should help improve the selection power as demonstrated in the Gaussian case. Simulation experiments validate the selection power improvement compared with the sequential construction in Section 4.2. Moreover, the bootstrap+ construction only requires $\mathbb{P}(A=e_{i}|A\neq 0_{m})$, which can be directly computed from the known empirical distribution $P_{\mathbf{X}}$ of the random vector $X$. This helps lower the computational cost, while avoiding the theoretical FDR inflation issues introduced from the estimation error on the distribution of $X$ at the same time.
2.3 Computation of $W$-statistics
With the above construction, we are now ready to compute the routine $W$-statistics similarly as in (Barber and Candès, 2015, Candes et al., 2018) through
$W_{i}=w_{i}\left\{(\widetilde{X}^{A},A,\widetilde{A}),Y\right\},$ (32)
for $i\in\{1,2,\cdots,m\}$. Similarly as in (Barber and Candès, 2015, Candes et al., 2018), the $W$-statistics needs to satisfy the flip-sign property on $A$ and $\widetilde{A}$,
$w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\}=\begin{cases}w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\},&j\notin S,\\ -w_{j}\left\{(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)},Y\right\},&j\in S.\end{cases}$ (33)
Here, $S$ represents any subset of $\mathcal{H}^{\gamma}_{0}$, and $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}$ denotes the operation of swapping the columns of $A$ and $\widetilde{A}$, while keeping $\widetilde{X}^{A}$ unchanged.
There are many approaches as shown in (Barber and Candès, 2015, Candes et al., 2018) in constructing such $W$-statistics satisfying the above property. In this paper, we adopt the following specific construction proposed by (Cao et al., 2024a) that searches in the relaxed neighborhood of the subspace constraint to optimize the selection power. Consider the following optimization problem for hyper-parameters $\lambda>0$, $\nu>0$:
$(\beta(\lambda),\gamma(\lambda),\widetilde{\gamma}(\lambda))=\mathop{\rm argmin}_{\beta,\gamma,\widetilde{\gamma}}$ $L(\widetilde{\mathbf{X}}^{A}\beta+\mathbf{A}\gamma+\widetilde{\mathbf{A}}\widetilde{\gamma},y)+\cdots$
$\cdots+\frac{1}{\nu}(\|D\beta-\gamma\|_{2}^{2}+\|D\beta-\widetilde{\gamma}\|_{2}^{2})+\lambda(\|\gamma\|_{1}+\|\widetilde{\gamma}\|_{1}),$ (34)
where the loss function $L$ is related with the model or link function, e.g., mean squared loss for linear regression, and cross-entropy loss for logistic regression. The $W$-statistics is constructed as
$W=|\gamma(\lambda)|-|\widetilde{\gamma}(\lambda)|.$ (35)
The hyper-parameters $\lambda>0$ and $\nu>0$ can be optimized to maximize the selection power. Here, $\lambda>0$ is the well-known $\ell_{1}$ penalization parameter. Meanwhile, $\nu>0$ is the parameter which controls the Euclidean relaxation gap between $D\beta$ and $\gamma$, and is previously introduced in Split LASSO (Cao et al., 2024a) in transformational sparsity problems. The effect of the hyperparameter $\nu$ for improving the selection power will be discussed by simulation experiments in Supplementary Material Section G.
With the above $W$-statistics, we are now ready to give the selectors of Model-X Split Knockoff. Let $q$ be our target FDR, two data-dependent threshold rules on a pre-set nominal level $q$ are defined as
[…]
or $T_{q},T_{q}^{+}=+\infty$ if the respective set is empty, where $\mathcal{W}=\{|W_{j}|:j=1,2,\cdots,m\}\backslash\{0\}$. The selectors for the Model-X Split Knockoff (+) are respectively defined as
$\widehat{\mathcal{S}}=\{i:W_{i}\geq T_{q}\}\mathrm{\ or\ }\widehat{\mathcal{S}}=\{i:W_{i}\geq T_{q}^{+}\}.$
2.4 Connections with Model-X Knockoff
The Model-X Split Knockoff is a generalization of the Model-X Knockoff method into the problem of transformational sparsity, whose application includes but not limited to pairwise comparisons, wavelet transforms, fused LASSO, and trend filtering. In the following, we argue that the Model-X Split Knockoff can achieve better selection power compared with the Model-X Knockoff in the case where both methods are applicable.
To be specific, consider the case of $D=I_{p}$, the model (1) degenerates to the traditional variable selection problem that Model-X Knockoff can deal with. In this case, we show by the following proposition that our method achieves at least the same selection power as the Model-X Knockoff.
Proposition 2.1.
Consider the case where $D=I_{p}$ in the model (1), there exists a pair of $(A,\widetilde{A})$ in construction of the Model-X Split Knockoff, such that the selection power of such a construction achieves at least the same selection power as the Model-X Knockoff.
Proof.
Take $A=X$ in the construction of the Model-X Split Knockoff, then $\widetilde{X}^{A}=X-AI_{p}=0$. Under this condition, Model (3) and the exchangeability condition (4) collapse precisely into the Model-X Knockoff framework. While this particular construction therefore achieves equivalent selection power, alternative constructions may offer superior performance. ∎
As shown in Section 2.2, taking $A=X$ might not be the optimal choice in the Model-X Split Knockoff in terms of the selection power, that our specific choice in Section 2.2 may achieve better selection power. Such a claim is further validated by simulation experiments in Section 4.1 and the Alzheimer’s disease study in Section 5.
3 FDR Control of Model-X Split Knockoff
In this section, theoretical guarantees for FDR control under the proposed Model-X Split Knockoff framework are provided. We start from the case where the distribution of $X$ is known and establish the exact FDR control. Then we apply the “leave one out” technique introduced in Barber et al. (2020) to establish robust FDR control when the distribution of $X$ can be estimated with a bounded deviation on KL-divergence.
3.1 Exact Model-X Split Knockoff
Now we are ready to present the false discovery rate control theorem of Model-X Split Knockoff when the distribution $X$ is known or bootstrapped exactly. In this setting, the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ exactly satisfies the exchangeability (4) and exact false discovery rate control is achieved.
Theorem 3.1 (FDR Control of Exact Model-X Split Knockoffs).
For any preset nominal FDR level $10<q<$ 1, if the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ satisfies the excangeability (4), the following statements holds for Model-X Split Knockoffs:
  1. 1.
    for the Model-X Split Knockoff, there holds
    $\mathrm{mFDR}=\mathbb{E}\left(\frac{|\{i\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|+1/q}\right)\leq q,$ (36)
  2. 2.
    for the Model-X Split Knockoff+, there holds
    $\mathrm{FDR}=\mathbb{E}\left(\frac{|\{i\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|\lor 1}\right)\leq q,$ (37)
where the $\mathrm{mFDR}$ represents the modified false discovery rate as introduced in Barber and Candès (2015) which adds $1/q$ in the denominator and should not have much impact when $|\widehat{\mathcal{S}}|$ is large.
Remark.
Theorem 3.1 will be a direct corollary of the robust FDR control Theorem 3.2 by setting the estimation error of $P_{X}^{*}$ to 0. The details will be discussed in Theorem 3.2, and the proof of Theorem 3.1 will therefore be omitted.
Example: Pairwise Comparisons.
Although the requirement that $(\widetilde{X}^{A},A,\widetilde{A})$ satisfies the excangeability (4) is generally unachievable in practice, there are some situations where such a condition reasonably holds.
For instance, consider the canonical example of pairwise comparisons in Section 2.2.2. In that case, the set of random vectors $(\widetilde{X}^{A}_{r},A,\widetilde{A})$ constructed by the bootstrap+ construction is based on bootstrapping (augmented with zeros) precisely the empirical distribution $P_{\mathbf{X}}$ of the random vector $X$. Therefore, it is straightforward to verify that $(\widetilde{X}^{A}_{r},A,\widetilde{A})$ satisfies the exchangeability (4), and exact false discovery rate control is achieved without estimating the distribution of $X$.
3.2 Approximate Model-X Split Knockoff
In practice, there are indeed many scenarios where the distribution of $X$ is unknown and needs to be estimated. In such cases, the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ only satisfies the exchangeability (4) approximately. To deal with such cases, we apply the “leave one out” technique introduced in Barber et al. (2020) to show that our method is robust to the estimation error on the distribution of $X$.
[…]
$\widehat{\mathrm{KL}}_{j}:=\sum_{i=1}^{n}\log{\left[\frac{P^{*}_{X}(\widetilde{\mathbf{X}}^{A}_{i}+\mathbf{A}_{i}D)}{P_{X}(\widetilde{\mathbf{X}}^{A}_{i}+\mathbf{A}_{i}D)}\frac{P_{X}\left\{\widetilde{\mathbf{X}}^{A}_{i}+\mathbf{A}_{i}(j)D\right\}}{P^{*}_{X}\left\{\widetilde{\mathbf{X}}^{A}_{i}+\mathbf{A}_{i}(j)D\right\}}\right]},$ (38)
where $\widetilde{\mathbf{X}}^{A}_{i}$, $\mathbf{A}_{i}$, $\widetilde{\mathbf{A}}_{i}$ represent the $i$-th row of $\widetilde{\mathbf{X}}^{A}$, $\mathbf{A}$, $\widetilde{\mathbf{A}}$ receptively, and $\mathbf{A}_{i}(j)$ represents the vector obtained by substituting the $j$-th element of $\mathbf{A}_{i}$ with the $j$-th element of $\widetilde{\mathbf{A}}_{i}$.
Clearly, if $P_{X}=P^{*}_{X}$, e.g. the distribution of $X$ is known, the KL divergence $\widehat{\mathrm{KL}}_{j}=0$ for $j\in\{1,2,\cdots,m\}$. Meanwhile, in the cases where the true distribution $P^{*}_{X}$ is unknown and is estimated by $P_{X}$, $\widehat{\mathrm{KL}}_{j}$ measures the impact of the estimation error to the $j$-th feature based on the observations $\widetilde{\mathbf{X}}^{A},\mathbf{A},\widetilde{\mathbf{A}}$.
Now we are ready to state the main theorem on the robust false discovery control of the approximate Model-X Split Knockoffs.
Theorem 3.2 (Robust FDR control of Approximate Model-X Split Knockoffs).
Suppose the set of random vectors $(\widetilde{X}^{A},A,\widetilde{A})$ satisfy the exchangeability (4) when $X\sim P_{X}$. Then when $X\sim P^{*}_{X}$, for any $10<q<$ 1 and $\epsilon\geq 0$, there holds:
  1. 1.
    for the Model-X Split Knockoff, there holds
    $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq q\cdot e^{\epsilon},$
    which leads to the following bound for modified FDR,
    $\mathrm{mFDR}=\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|+q^{-1}}\right)\leq\min_{\epsilon\geq 0}\left\{q\cdot e^{\epsilon}+\mathbb{P}\left(\max_{j\in\mathcal{H}_{0}^{\gamma}}\widehat{\mathrm{KL}}_{j}>\epsilon\right)\right\},$
  2. 2.
    for the Model-X Split Knockoff+, there holds
    $\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\text{ and }\widehat{\mathrm{KL}}_{j}\leq\epsilon\}|}{|\widehat{\mathcal{S}}|\vee 1}\right)\leq q\cdot e^{\epsilon},$
    which leads to the following bound for FDR,
    $\mathrm{FDR}=\mathbb{E}\left(\frac{|\{j:j\in\widehat{\mathcal{S}}\cap\mathcal{H}_{0}^{\gamma}\}|}{|\widehat{\mathcal{S}}|\vee 1}\right)\leq\min_{\epsilon\geq 0}\left\{q\cdot e^{\epsilon}+\mathbb{P}\left(\max_{j\in\mathcal{H}_{0}^{\gamma}}\widehat{\mathrm{KL}}_{j}>\epsilon\right)\right\}.$
The proof of Theorem 3.2 will be presented in Supplementary Material Section B.
Remark.
For the case that $P_{X}=P^{*}_{X}$, taking $\epsilon=0$ in Theorem 3.2 directly leads to Theorem 3.1. In this sense Theorem 3.1 can be regarded as a corollary of Theorem 3.2.
Theorem 3.2 states that the loss in the FDR control vanishes when the estimation on the distribution of $X$ becomes accurate. In the following, we will give a concrete example example on the robust FDR control when the distribution of $X$ is a normal distribution.
Example: Normal Distributions
In this following, we provide concrete computation of $\max_{j}\widehat{\mathrm{KL}}_{j}$ when $X$ follows the normal distribution $\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. In this example, we construct the Model-X Split Knockoff in the way presented by Equation (25) and (30).
Theorem 3.3.
Denote $\Theta:=\Sigma^{-1}_{X}$ and $\widehat{\Theta}:=\widehat{\Sigma}^{-1}_{X}$, where $\widehat{\Sigma}_{X}$ is an estimation of $\Sigma_{X}$. Suppose $X\sim\mathcal{N}(\mathbf{0}_{p},\Sigma_{X})$. Let $\Lambda_{X}$ the maximum eigenvalue of $\Sigma_{X}$
$\delta_{\Theta}$ $:=4\alpha\|D_{j}\|_{2}\|\Delta\|_{2}\left(\sqrt{2p\Lambda_{X}}+2\alpha\|D_{j}\|_{2}^{2}\right)$
for any choice of $\alpha>0$ such that $\widehat{\Sigma}_{X}-\alpha D^{T}D\succeq 0$ and $A\sim\mathcal{N}(0,\alpha I_{m})$. If $\frac{\log{(m+p)}}{n}=o(1)$, with probability at least $1-\frac{2}{p}-\frac{2}{m}$, there holds
$\max_{j}\widehat{\mathrm{KL}}_{j}\leq 2\delta_{\Theta}\sqrt{n\log{m}}\{1+o_{p}(1)\}.$
Remark.
In many scenarios including total variation or multiple comparisons, $\|D_{j}\|_{2}$ is at constant levels, so we only need to focus on the estimation error of $\Theta$ to establish bounds on $\delta_{\Theta}$ and $\max_{j}\widehat{\mathrm{KL}}_{j}$. To be specific, the estimation error $\|\Theta-\widehat{\Theta}\|_{2}$ needs to be $o_{p}\left(\frac{1}{\sqrt{n\log p}}\right)$ in order to make $\max_{j}\widehat{\mathrm{KL}}_{j}$ sufficiently small.
To achieve this goal, unlabeled data can be utilized for estimation of $\Theta$, which are typically more accessible and thus considerably outnumber labeled data. To be specific, consider the case where there are $N\gg n$ unlabeled samples. Given that $X$ follows a multivariate Gaussian distribution, a simple sample covariance matrix $S_{N}:=\frac{1}{N}(X^{u})^{\top}X^{u}$ computed over unlabeled data $X^{u}$ can be used as the estimation, which has $\|\Theta-\widehat{\Theta}\|_{2}\sim O_{p}\left(\sqrt{\frac{p\log{p}}{N}}\right)$. In this case, $\max_{j}\widehat{\mathrm{KL}}_{j}\sim o_{p}(1)$ is obtained as long as
$2\delta_{\Theta}\sqrt{n\log{m}}\asymp\sqrt{\frac{p\log{p}}{N}}\sqrt{pn\log{m}}=o_{p}(1),$
i.e. if $N\gg np^{2}\log^{2}{(m+p)}$. Further, if the inverse covariance matrix is sparse, one can employ the Graphical Lasso for estimation:
$\widehat{\Theta}_{\lambda}=\arg\min_{\Theta\succeq 0}\left\{-\log\operatorname{det}(\Theta)+\left\langle\Theta,S_{N}\right\rangle+\lambda\sum_{j\neq k}\left|\Theta_{jk}\right|\right\}.$
Under certain conditions, Ravikumar et al. (2011) shows that $\|\Theta-\widehat{\Theta}\|_{2}\asymp\sqrt{\frac{d^{2}\log{p}}{N}}$, where $d$ denotes the maximum degree. This property only requires $N\gg npd^{2}\log^{2}{p}$ to make $\max_{j}\widehat{\mathrm{KL}}_{j}$ to be $o_{p}(1)$.
4 Simulation Experiments
In this section, the performance of the proposed Model-X Split Knockoff method is evaluates in two canonical settings, the normal distribution setting and the pairwise comparisons setting. The simulation results validates that the Split Knockoff method achieves desired false discovery rate control as well as relatively high selection power, specifically higher compared with Model-X Knockoff when both are applicable.
4.1 Normal Distribuions
Experimental Setting: In this section, simulation experiments are conducted under the case where the distribution of the design $X$ follows the normal distribution. In particular, rows of the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is generated independent and identically distributed (i.i.d.) from a multivariate normal distribution $\mathcal{N}(0_{p},{\bm{\Sigma}})$, where the covariance matrix ${\bm{\Sigma}}$ has unit diagonal entries (${\bm{\Sigma}}_{i,i}=1$) and correlation structure ${\bm{\Sigma}}_{i,j}=c^{|i-j|}$ for $i\neq j$ with correlation parameter $c=0.5$:
${\bm{\Sigma}}=\begin{pmatrix}1&c&c^{2}&c^{3}&\cdots&c^{p-1}\\ c&1&c&c^{2}&\cdots&c^{p-2}\\ c^{2}&c&1&c&\cdots&c^{p-3}\\ c^{3}&c^{2}&c&1&\cdots&c^{p-4}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ c^{p-1}&c^{p-2}&c^{p-3}&c^{p-4}&\cdots&1\end{pmatrix}.$
The true coefficient vector ${\beta}^{*}\in\mathbb{R}^{p}$ is defined by
$\beta^{*}_{i}=\begin{cases}A&i\leq k,\quad i\equiv 0,-1\pmod{3},\\ 0&\text{otherwise}.\end{cases}$ (39)
In this section, we consider the logistic regression, where the response $\mathbf{Y}\in\mathbb{R}^{n}$ is generated from
$\mathbf{Y}\sim\text{Bernoulli}\left\{\sigma(\mathbf{X}\beta^{*})\right\}.$ (40)
Here, $\sigma(x)$ denotes the sigmoid function, i.e. $\sigma(x)=\frac{1}{1+\exp(-x)}$.
The transformational sparsity is specified by the linear transformation matrix $D\in\mathbb{R}^{m\times p}$ such that $\gamma^{*}=D\beta^{*}$ is sparse. Based on the above specific choice of $\beta^{*}$, the following three kinds of $D$ is considered in this simulation:
  • $\beta^{*}$ itself is sparse, so that one can take $D_{1}=I_{p}$, where $I_{p}$ is the identity matrix of size $p\times p$, leading to $m=p$.
  • $\beta^{*}$ is a uni-dimensional piecewise constant function, so that one can take $D_{2}$ as the 1-D graph difference operator on a line. Specifically, $D_{2}\in\mathbb{R}^{(p-1)\times p}$, where $D_{2}(i,i)=1$, $D_{2}(i,i+1)=-1$ for $i\in\{1,\dots,p-1\}$, and $D_{2}(i,j)=0$ for other pairs of $(i,j)$. In this case, $m=p-1<p$.
  • Combining the above two cases, one can take $D_{3}=\begin{pmatrix}D_{1}\\ D_{2}\end{pmatrix}\in\mathbb{R}^{(2p-1)\times p}$, where $m=2p-1>p$.
For the case where $D_{1}=I_{p}$, the performance of Model-X Knockoffs is also evaluated for comparisons. Throughout this section, the target FDR level $q$ is set to be $0.2$, while the coefficients in the above simulation design are taken to be $p=100$, $k=20$, and $A=1$. The Model-X Knockoff method is implemented using the official repository [^4] https://github.com/msesia/knockoff-filter.
The Model-X (Split) Knockoff copies are generated from the data matrix $\mathbf{X}$ through the estimated covariance matrix $\widehat{\bm{\Sigma}}=\mathbf{X}^{T}\mathbf{X}/n$. Specifically, the Split Knockoff copies $A$ and $\widetilde{A}$ are generated through Equation (25) and (20). To select the tuning parameters $\lambda$ and $\nu$ in Equation (2.3), cross-validation is applied to identify the optimal values, $\widehat{\lambda}$ and $\widehat{\nu}$, based on the test loss. To solve the optimization problem (2.3), we leverage the Embedded Conic Solver (ECOS) algorithm for convex optimization with the CVX [^5] https://www.cvxpy.org package.
4.1.1 Comparisons with Model-X Knockoffs when $D=I_{p}$
(i) Performance of Model-X Knockoff
(ii) Performance of Model-X Split Knockoff
Figure 1: Performance of Model-X Knockoffs and Model-X Split Knockoffs when the number of samples varies in the case $D=D_{1}=I_{p}$. The logistic regression is implemented for $q=0.2$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
In Figure 1, we compare the performance of Model-X Knockoffs and Model-X Split Knockoffs when $D=D_{1}=I_{p}$, i.e. the case where the transformation sparsity problem is degenerated into the classical direct sparsity problem. In both figures, with the increase of sample size, the selection power of both methods gradually increases from 0 to 1, while the FDR of both methods are under control for all sample sizes.
In terms of selection power, the Model-X Split Knockoff demonstrably yields superior results across all tested sample sizes, a finding supported by Proposition 2.1. We hypothesize that this performance gain is a direct consequence of the improved incoherence conditions, which arise from the orthogonal design of $A$ and $\widetilde{A}$ in Equations (25) and (30).
4.1.2 Performance of Model-X Split Knockoffs on Non-trivial $D$
(i) $D=D_{2}$
(ii) $D=D_{3}$
Figure 2: Performance of Model-X Split Knockoff when the number of samples varies in the cases where $D=D_{2}$ or $D=D_{3}$. The logistic regression is implemented for $q=0.2$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
In Figure 2, the performance of Model-X Split Knockoffs is evaluated in the scenarios where $D$ is non-trivial, i.e. $D\neq I_{p}$. In both cases, the FDR of Model-X Split Knockoffs is under control universally for all choices of sample sizes. Meanwhile, the selection power of Model-X Split Knockoffs follows a similar trend as it is in the case where $D=I_{p}$, i.e. the selection power gradually increases from 0 to 1 when the sample size enlarges.
4.2 Pairwise Comparisons
Experimental Settings: In this section, the pairwise comparison problem among $p$ objects with $n$ noisy observations of comparisons is considered. The scores for $p$ objects, denoted as $\beta^{*}\in\mathbb{R}^{p}$, are generated from a standard normal distribution $\mathcal{N}(0,1)$. To introduce sparsity into the dataset, the scores in $\beta^{*}$ are truncated according to varying sparsity ratios $k$ set at 0.1, 0.3, and 0.5. Specifically, only the first $k\cdot p$ components of $\beta^{*}$ are retained, while the remaining components are set to zero.
The rows of the pairwise comparison design matrix $\mathbf{X}\in\{0,1,-1\}^{n\times p}$ is generated independently in the following way. For the $i$-th row, two indices $1\leq j<k\leq p$ are chosen uniformly at random with probability $\frac{2}{p(p-1)}$ for each possible pair of $(j,k)$. Then we set $\mathbf{X}_{i,j}=1$, $\mathbf{X}_{i,k}=-1$, and $\mathbf{X}_{i,l}=0$ for $l\neq j,k$. This specific construction ensures that each row of $\mathbf{X}$ represents a (noisy) comparison between two objects.
In this simulation experiment, the performance of Model-X Split Knockoff is evaluated under the Bradley-Terry model, where the response vector $\mathbf{Y}\in\mathbb{R}^{n}$ is generated by Equation (40). The transformational matrix $D\in\mathbb{R}^{\frac{p(p-1)}{2}\times p}$ is taken to be the graph difference operator on a full connected graph with $p$ vertices to conduct pairwise comparisons.
(i) Bootstrap+
(ii) Sequential
Figure 3: Performance of Model-X Split Knockoff as the Number of Features Varies for Bootstrap+ and Sequential Constructions. Simulation experiments are conducted using the Bradley-Terry Model of pairwise comparison with $q=0.2$ and a sparsity level of $0.5$ for $\beta^{*}$ (i.e., $k/p$ in Equation (39)). The curves represent average performance, while the shaded areas indicate 80% confidence intervals based on 200 simulation runs.
Figure 3 compares the performance of bootstrap+ constructions of Model-X Split Knockoff copies against the sequential constructions. It shows that, although the two types of constructions achieve desired FDR control, the selection power of the bootstrap+ constructions is higher compared with the sequential constructions when the number of features varies. This is explained in Section 2.2.2 that the bootstrap+ constructions forces $\widetilde{X}^{A}$ to be independent from $A$, which helps improve the incoherence conditions and thus the selection power.
Figure 4: Runtime(Second) of Model-X Split Knockoff for bootstrap+ and sequential methods. Results are averaged on 200 simulation runs with the number of samples $n=1000$, and the number of features varying from 5 to 30.
Moreover, when $p$ is large, the sequential construction may need to estimate too many conditional distributions, that the computation cost explodes. On the other hand, the bootstrap+ constructions does not suffer from this issue, and handles such a case easily. As shown in Figure 4, the sequential method exhibits a much steeper increase in running time compared to the bootstrap+ method, which requires significantly less computation time under the same settings.
Due to the above two points, the bootstrap+ construction will be the default construction to implement Model-X Split Knockoff for pairwise comparisons in the following sections.
(i) Sparsity Level $k=0.1$
(ii) Sparsity Level $k=0.3$
(iii) Sparsity Level $k=0.5$
Figure 5: Performance of Model-X Split Knockoff when the number of samples varies. Simulations experiments are implemented through Bradley-Terry Model of pairwise comparison for $q=0.2$ with different sparsity level of $\beta^{*}$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
Figure 5 illustrate the performance of Model-X Split Knockoffs when the number of samples and the sparsity level varies. In this figure, we adopt the bootstrap+ constructions of the Model-X Split Knockoff copies. Again, for all cases, the FDR of Model-X Split Knockoffs is under control. For each sparsity level, the selection power of Model-X Split Knockoffs gradually increases when the sample size increases.
5 Alzheimer’s Disease
We employ the Model-X Split Knockoff method in this study to analyze lesion regions and their connectivity within the brains of individuals affected by Alzheimer’s Disease (AD), which has been the subject of growing research focus.
5.1 Dataset Description
ADNI  [^6] http://adni.loni.ucla.edu dataset is collected by the structural Magnetic Resonance Imaging (MRI), which consists of 752 samples divided into three types: 126 AD, 433 Mild Cognitive Impairment(MCI), and 193 Normal Controls (NC). Each image is pre-processed by the toolbox Statistical Parametric Mapping (SPM) for segmentation into gray matter(GM), white matter (WM), and cerebral spinal fluid (CSF), followed by Automatic Anatomical Labeling(AAL) to partition the whole brain into 90 regions with the volume of each region.
To utilize the dataset, we construct a design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n=752$ and $p=90$, representing the volumes of $90$ processed brain regions for all $752$ samples. Specifically, $\mathbf{X}_{i,j}$ denotes the column-wise normalized volume of region $j$ for sample $i$. In this section, $\mathbf{Y}\in\mathbb{R}^{n}$ is taken as the Alzheimer’s Disease Assessment Scale (ADAS), a measure initially developed to evaluate the severity of cognitive dysfunction (Rosen et al., 1984), which was later shown to effectively distinguish between clinically diagnosed Alzheimer’s disease and normal controls (Zec et al., 1992).
5.2 Region Selection
Table 1: Selected Regions by Model-X Knockoff and Model-X Split Knockoff on Alzheimer’s Disease ($q=0.2$). The full names of the regions are provided in Supplementary Material Section I.
Region Model-X Knockoff Model-X Split Knockoff
Frontal Inf Orb L
Frontal Inf Oper L
Parietal Inf R
Frontal Mid R
Rolandic Oper L
Rolandic Oper R
Cingulum Post L
Cingulum Post R
Hippocampus L
Hippocampus R
Lingual L
Precuneus R
Fusiform R
Supramarginal R
Temporal Mid L
Temporal Mid R
The region selection results of Model-X Split Knockoff and Model-X Knockoff are shown in Table 1. For Model-X Split Knockoff, the $W$ statistics is computed based on the cross-validated $\lambda$ and $\nu$ on a grid of $\log{\lambda}$ between $-2$ and $-1$ with a step size of $0.05$ and a grid of $\log{\lambda}$ between 00 and $1$ with a step size of $0.05$.
Overall, Model-X Split Knockoff identifies 15 brain regions associated with Alzheimer’s disease, whereas the standard Model-X Knockoff selects only 8, demonstrating the enhanced selection power of our approach.
Both methods select the hippocampus, middle temporal lobes (”Temporal Mid”), middle frontal gyrus (”Frontal Mid R”), and several cerebral cortex areas such as the lingual gyrus and fusiform gyrus. Most of these regions have been previously reported as potential biomarkers for Alzheimer’s disease in the literature (Lee et al., 2020, Peters et al., 2009, Yang et al., 2019). Specifically, the hippocampus and middle temporal lobes are known to be involved in memory function (Vemuri and Jack, 2010) and are well-established biomarkers in the progression of Alzheimer’s disease (Lee et al., 2020). Additionally, Yang et al. (2019) reported a significant reduction in cortical thickness in AD patients compared to normal controls (NC) across cerebral cortical regions, including the lingual gyrus and fusiform gyrus. Besides, Busatto et al. (2008) found that some sub-regions in the middle frontal gyrus suffered from degeneration in the early stage of the disease. These regions may exhibit strong discrimination in our dataset, causing our methods to favor this selection.
Besides, most of the regions that are additionally selected by our method have also been reported to undergo early degeneration in previous studies (Van Hoesen et al., 2000, Cajanus et al., 2019, Pengas et al., 2010, Lee et al., 2020). In particular, the inferior frontal gyrus plays multiple roles, including speech and language processing (Van Hoesen et al., 2000, Cajanus et al., 2019). Additionally, significant structural changes in nodal centrality within the rolandic operculum have been observed in the AD population (Yao et al., 2010, Bajo et al., 2015). Besides, the posterior cingulate cortex, which is involved in episodic memory, was found to exhibit atrophy at the earliest clinical stages of sporadic AD (Pengas et al., 2010, Lee et al., 2020).
5.3 Connection Selection
In this section, we detect abnormal connections between brain regions. We represent brain connectivity as a graph $G=(V,E)$, where $V$ denotes the brain regions and $E$ represents edges connecting adjacent regions. Based on this formulation, the transformational matrix $D$ is defined as a graph difference operator, such that $D(\beta)(i,j)=\beta_{i}-\beta_{j}$ for adjacent brain regions $(i,j)$. Since adjacent brain regions typically show similar activity, a significant signal on an edge suggests connections between relatively stable regions and severely atrophied regions contributing to the disease.
Table 2: Selected Connections by Model-X Split Knockoff on Alzheimer’s Disease ($q=0.2$). The full names of the regions are provided in Supplementary Material Section I.
Region 1 Region 2 Model-X Split Knockoff
Hippocampus L Cingulum Post L
Hippocampus L Fusiform L
Hippocampus L Precuneus L
Hippocampus L Putamen L
Hippocampus L Thalamus L
Hippocampus L Heschl L
Hippocampus R Lingual R
Hippocampus R Putamen R
Hippocampus R Temporal Sup R
Hippocampus R Temporal Inf R
Hippocampus R Insula R
Amygdala L Frontal Inf Orb L
Amygdala R Temporal Pole Sup R
Lingual L Fusiform L
Lingual L Thalamus L
Lingual L Cingulum Post L
Lingual R Occipital Mid R
Lingual R Fusiform R
Lingual R ParaHippocampal R
The abnormal connections identified by our method are presented in Table 2. Notably, many of the detected connections involve the hippocampus, which exhibits the earliest and most significant atrophy in Alzheimer’s disease (AD) (Greicius et al., 2003, Yang et al., 2019, Cajanus et al., 2019, Yao et al., 2010).
In addition, we identify two selected connections involving the amygdala. Anatomically adjacent to the hippocampus, the amygdala is also part of the medial temporal lobe and plays a key role in memory processing (Li et al., 2023). Notably, Nelson et al. (2018) reported that the amygdala is a primary site of amyloid plaque deposition in the early stages of AD—a hallmark that has been widely hypothesized as a defining feature of AD. Consistently, the amygdala has been shown to undergo early degeneration during disease progression (Ferry, 2012, Stouffer et al., 2023).
Moreover, the lingual regions were frequently involved in the selected connections. This finding is consistent with previous studies showing that the lingual gyrus—associated with visual processing and visual memory—plays an important role in the early detection of Alzheimer’s disease. Notably, impairments in lingual function have been suggested as another early indicator of AD (Yang et al., 2019, Van Hoesen et al., 2000).
6 Pairwise Comparison: WorldCollege Dataset
In this section, we apply our method to the World College dataset [^7] http://www.allourideas.org/worldcollege. This dataset contains 9,408 user-annotated pairwise comparisons among 261 colleges, collected via the All Our Ideas crowdsourcing platform; 8,544 of these comparisons were deemed valid. A diverse group of 400 annotators from various countries were asked to choose which university in a given pair they found more attractive to attend.
For our experiments, we subsample the dataset to exclude colleges with very few annotated comparisons. Our specific sub-dataset comprises $p=69$ colleges that rank in the top 80 of the QS rankings among the original 261 colleges. This subsampling yields $n=2112$ annotated pairwise comparisons among these 69 colleges, resulting in $m=608$ unique pairwise comparisons. More details on the QS rankings are in Supplementary Material Section I.
In this experiment, the Bradley-Terry model is adopted, and thus the transformation matrix $D\in\mathbb{R}^{m\times p}$ is taken as the graph difference operator of the subsampling graph, i.e. for each pair of colleges ($i_{1}<i_{2}$) compared by the annotator, there exists an unique row $j$ in $D$, such that $D_{j,i_{1}}=1$, $D_{j,i_{2}}=-1$, and $D_{j,k}=0$ for $k\neq i_{1},i_{2}$.
Figure 6: Selected Pairs of Colleges with Significant Preference Differences. Each point in the plot represent the pair of colleges with respective QS ranks marked by the column and row indices. To be specific, the dark blue points represent the pairs of colleges selected by our method to have significant preference differences, while the light blue points represent the pairs of colleges that are included in $D$. The diagonals are marked in deeper dark blue for better visualization.
Figure 6 presents the selection results for pairs of colleges with significant preference differences. As the figure shows, our method tends to select pairs located further from the diagonal with higher frequency than those closer to it. This indicates that our method effectively identifies university pairs with substantial QS ranking differences, which are more likely to represent true discoveries. For instance, our method identified the pairs Massachusetts Institute of Technology (MIT) (QS rank 1) and Osaka University (QS rank 63), and MIT (1) and Tohoku University (75), both exhibiting large ranking gaps. Other examples include Stanford University (2) and Lund University (73), and Harvard University (3) and Korea Advanced Institute of Science and Technology (47). These findings support the effectiveness of our approach in uncovering university pairs with notable differences in global academic standing.
Figure 7: Selection Frequency of Pairwise Comparisons by Model-X Split Knockoffs Across 100 Runs. The bar chart displays the average selection frequency for college pairs identified by our method. Each bar corresponds to pairs within $D$ whose QS ranking differences are categorized by the range shown on the $X$-axis.
To further illustrate our method’s selection tendencies, Figure 7 presents a frequency plot displaying the selection frequencies of pairs against their QS ranking differences. The figure clearly shows an increasing trend in selection frequency as the QS ranking difference between pairs grows. This further validates our claim that our method tends to select university pairs with significant QS ranking differences and likely maintains a low False Discovery Rate (FDR). More specifically, our method consistently detects connections between colleges exhibiting substantial differences in their global rankings. For instance, connections between the Massachusetts Institute of Technology (rank 1) and Osaka University (rank 63), or Harvard University (rank 3) and Tohoku University (rank 75), were detected in 73 and 89 out of 100 experiments, respectively. The detected connections frequently involve pairs with rank differences exceeding 60, demonstrating our method’s particular sensitivity and effectiveness in uncovering relationships between institutions with large rank differences.
7 Conclusion
In this paper, we introduce the Model-X Split Knockoff method, an extension of both the Split Knockoff and Model-X Knockoff frameworks. This method effectively controls the false discovery rate in transformation selection problems across a broad range of non-linear statistical models. To overcome the inherent challenge posed by models combining a random design matrix with a deterministic linear transformation, we introduce a novel auxiliary randomized design. This randomization interacts with both the random design and the deterministic transformation, enabling the extraction of information related to transformational sparsity and successfully bypassing associated difficulties. Moreover, the flexibility of this auxiliary randomized design allows Model-X Split Knockoffs to achieve superior selection power compared to standard Model-X Knockoffs when both are applicable. Simulation experiments and real-world applications to Alzheimer’s Disease and the World College ranking validate the proposed method’s effectiveness.